OpenStudy (anonymous):

Consider the solid S described below. The base of S is the triangular region wi

OpenStudy (anonymous):

Consider the solid S described below. The base of S is the triangular region with vertices (0, 0), (2, 0), and (0, 2). Cross-sections perpendicular to the x-axis are equilateral triangles. Find the volume V of this solid.

OpenStudy (anonymous):

Are you kidding me? I appreciate you help but that is like asking me something in chinese.

OpenStudy (dumbcow):

haha :)

OpenStudy (anonymous):

i dont get this question

OpenStudy (dumbcow):

it best to try to draw it out to get an idea of the visualization. The base is a right triangle and if you cut it vertically each cross-section is a equilateral triangle. volume is the infinite sum of all the cross-sectional areas

OpenStudy (dumbcow):

goal is to represent area of equilateral triangle in terms of x. then integrate from 0 to 2 with respect to x

OpenStudy (dumbcow):

Area of equilateral triangle = \[\frac{\sqrt{3}}{4}s^{2}\] where s is length of a side s is represented by the length from the x-axis to the hypotenuse of base for each cross-section. The line from (0,2) to (2,0) is \[y = 2-x\] so s = 2-x As x gets bigger the cross-sectional areas get smaller The integral should look like this: \[V = \int\limits_{0}^{2}\frac{\sqrt{3}}{4}(2-x)^{2} dx\]

OpenStudy (anonymous):

Are you serious? Thanks but I don't get it.

OpenStudy (dumbcow):

Finding volume of a solid is tough to teach online...you need a visual. this is a calc 2 type problem is this question for you dumbunny? or vaska?

OpenStudy (dumbcow):

so did that help? can you do the integration from here?

OpenStudy (anonymous):

yes, thanks...

OpenStudy (dumbcow):

ok your welcome in general for these type of problems \[V = \int\limits_{}^{}A(x) dx\] where A is area of cross-section